40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6679     Entry time: Thu May 24 19:39:18 2012
Author: Suresh 
Type: Update 
Category: IOO 
Subject: MC and WFS alignment adjusted 

[Yuta, Suresh]

We found that the MC was not locking and that the alignment between PSL and MC was too poor to obtain a TEM00 mode in the MC.   To correct the situation we went through the following steps:

1) We burt restored the MC alignment slider values to their values at 3:07 AM of today

2) We turned off the MC-autolocker and the ASC signal to the coils.   Then aligned the PSL beam into the MC (with the MC servo loop off) to obtain the TEM00 mode.  We had to adjust the zig-zag at the PSL output by quite a bit to maximise MC transmission.

3) We then centered the spot on the MC2 face and centered the transmitted beam on the MC2_TRANS_QPD

4) Next, we centered the beams on the MC_WFS sensors.

5) Turning on the WFS loops after this showed that everything works fine and WFS loops do not accumulate large offsets.

 

 

ELOG V3.1.3-